A small plane travels 200mph with the wind and 120 mph against it. Find the speed of the wind and the speed of the plane in still air?
Answer by checkley77(12844)

(200-120)/2=80/2=40 MPH. IS THE WIND SPEED.
200-40=160 IS THE SPEED OF THE PLANE IN STILL AIR.
12-+40=160 DITTO.